### v1

@@ 5.2 Common Event Attribution Tasks
As described in section 4, we ran the experiments on all 66 distinct monthly datasets we collected and processed from the GDELT project. The results of those experiments appear in Figures 7. To summarize these results, we also constructed the Following Table 3 which contains the average performance of all methods, across all datasets, measured metrics, and tasks. In Figure 7, we present the performance of each method on both tasks: daily and monthly common event attribution across all 66 monthly separated datasets of GDELT data. This plot enables further investigation of the benefits and shortcomings of each method, by presenting their relationship and behavior given a dataset.
Additionally, in Figures 6, we visualize the distribution of the performance of all methods in the context of metrics: ROC and PR AUC, and tasks: monthly and daily, across all the datasets. Moreover, we performed statistical analysis to determine whether there was a significant difference in the methods. We concluded that the concatenated approach provides better performance in both tasks compared to the standalone baseline and SiameseNet by the Precision-Recall AUC evaluation.
@@ 5.3 Statistical significance evaluation
We perform statistical analysis to verify the existence of significant differences among the evaluated methods. As suggested by [73], we use the Friedman test [69] on the results of the aforementioned approaches to identify which obtained better results regarding each designed task and metric - across all datasets.
The tests were used with a significance level of 5%. Specifically, the rank results established by the statistical the method provides the criterion for evaluating the approaches, considering that the higher the value of the calculated rank, the better the result will be in this approach. In our case, the null hypnosis was rejected in each context, as can be seen in Table 5. The rejection of the null hypothesis of the Friedman test indicates the existence of statistical differences between the methods but does not define which of them are statistically different. For that, the Nemenyi post-hoc test [70] was applied to make multiple comparisons using the values of the percentages of means for each approach evaluated in all the executed datasets in relation to their context, as can be seen in Table 6. The results of the statistical analysis are presented in A.4.
Therefore, in all cases, there were significant differences between the baseline SIF approach and the Concatenated approach. While by the Precision-Recall metric, there have also been significant differences between our proposed SiameseNet and the Concatenated approach.
@@ 6 Discussion
First, the utilization of entities, themes, and events to construct news embeddings has shown promising results, this is reflected in Tables 3, 1. Even though the task of shared event detection between pairs of articles in a said timeframe is challenging as discussed in the Introduction section, all proposed methods achieved extremely high ROC AUC scores, all above 0.95. While also providing impressive results with regard to the more challenging metric of PR AUC.
@@
Second, the utilization of the semi-supervised approach, which considers also cross-document latent information, given the tagged events appears to contribute to our method performance. From Figure 6 it is evident that the semi-supervised approach achieves significant improvement compared to the SIF vaseline by the ROC AUC metric on all tasks. Additionally from analyzing Figure 7, we see that even though there is a high correlation between the method’s performance, a variance in performance exists. Meaning we found some cases in which the baseline performs better than the semi-supervised approach. We attribute this behavior to two reasons: (a) we found a moderate correlation of 0.369 for each dataset size and the gap in performance between the SIF and semi-supervised approach, more data led to a bigger gap in favor of the Siamese network, (b) a lack of adaptation of the Siamese network to the given dataset or even task led to variances in its results.
@@
Third, our method which generates news article embeddings was able to produce embeddings for both articles and entities in the process, on a huge scale of news data, without the use of GPU in the process. Using the resources described in A, we ran every step of the pipeline, and every experiment on all the generated datasets in parallel. For each entity and theme we monitored, over 320,000 overall (see Table 2), we generated topical embeddings for every period they appeared in from 07/2015 to 01/2021. We have considered all articles from dozens of different monitored media sources and their affiliates, as presented in Figure 4, resulting in a dataset of over 850,000 articles. This enables researchers with fewer resources to investigate the large-scale of online news, using this method.
@@
Fourth, we constructed topical monthly entity embeddings for over 320,000 distinct entities and themes. The pooling mechanism that controlled the training set of the GloVe algorithm served as a critical point in this research. This mechanism not only allowed us to control the resource utilization of our suggested method but also fine-tuned the representation of each entity to a recent time frame as its articles’ publication date. In Table 4 we further explain and demonstrate real-world examples of the importance of time topical embeddings. We plan to share these topical entity embeddings with the research community. These topical embeddings could serve other studies that seek to examine how relationships between entities change over time.
Fifth, The experimental results, as presented in the previous section, demonstrate the superiority of our proposed method compared to the baseline method SIF and the Siamese network. This can be detected in both Figures 6, 7, and in Table 3. We detected only three instances out of 66 datasets and two tasks in which the SIF approach surpassed our technique. Across both daily and monthly event attribution tasks, our technique consistently outperformed SIF, surpassing it by 2.57% in PR-AUC and 1.78% in ROC-AUC for the daily task, and by 2.15% in PR-AUC and 1.67% in ROC-AUC for the monthly task. We did see, that in the less restrictive metric of ROC AUC, the semi-supervised approach and the concatenation approach received similar results. However, as brought from the statistical analysis, performed using the Friedman and Nemenyi tests with a 5% we validated the significant improvement of the concatenation approach on the PR AUC metric, on both tasks. As detailed in Section 5.3 and A.4.
@@
Lastly, we observed that in the cases in which a better performance is yielded by the Siamese network or SIF Methods compared to the Concatenated approach, there was a high gap between the SIF approach and the Siamese approach, leading to failure of the concatenated approach to successfully account for additional information by utilizing both approaches. Resulting in lesser performance than taking the best approach given the dataset.
@@
Finally, the methods and method presented in this research have some limitations worth noting: First, it is reliant on the historical existence of event tagging for news articles. Which might not be available, or reliable enough. Second, as we suggested by the results our methodology is rather static and the pipeline should preferably be manually tuned to provide adaptive embeddings to the required task, these concerns both in the semi-supervised training process and in the feature selection phase.
@@ 7 Conclusions and Future Work
This research has presented a novel approach for generating event-based embeddings for news articles, addressing the limitations of existing methods that rely solely on the textual content of documents, disregard runtime and memory costs, and do not rely on cross-document information. By leveraging the historical context of entities, themes, and their associations with specific events, our proposed method aimed to capture the latent relationships between these elements, ultimately enhancing the representation of news articles in a vector space.
We showed our method’s ability to produce articles’ embeddings that consider the article’s published time, represent the latent event information that appears in them, and do so in relatively low resource usage on large-scale datasets. We also demonstrated the method’s ability to surpass the arbitrary aggregation function of the baseline SIF approach. Moreover, we utilized the synthesis between the SIF approach and the semi-supervised Siamese network approach.
This research offers multiple directions for possible future work:
investigating the interpretability and explainability of the generated embeddings could be a valuable research direction. Understanding the factors that contribute to the embeddings’ representations and their relationships with events and entities could lead to more transparent and trustworthy models, as well as facilitate the incorporation of human expertise and domain knowledge into the embedding generation process.
This also paves the way for analyzing the relationships between certain entities and tense or biased media coverage. In the future, we may apply a holistic approach to event clustering based on the methods proposed in this work, to facilitate possible research in the media slant or political news.
@@ A.3 Embedding Generation Notes
Entity Entity Neighbour GloVe GloVe+Pool Mechanism
@@
Fith RS Alexander_Litvinenko

### v2

@@ 5.2 Common Event Attribution Tasks
As described in Section 4, we ran the experiments on all 66 distinct monthly datasets we collected and processed from the GDELT project. The results of those experiments on our suggested approach which considered both events and the time-based entity embeddings appear fully in Figure 8.
To summarize these results, we also constructed the following Table 3, which contains the average performance of all methods, across all datasets, measured metrics, and tasks.
Additionally, in Figures 6, we visualize the distribution of the performance of all methods in the context of metrics (ROC and PR AUC) and tasks, both monthly and daily, across all the datasets. Moreover, we performed statistical analysis to determine whether the methods had a significant difference. We concluded from the Precision-Recall AUC evaluation that the concatenated approach provides better performance in both tasks compared to the standalone baseline and Siamese network.
5.3 Comparison to Existing Method
The same experiments were performed on the spaCy large CPU model, the results for those experiments are shown fully in Figure 9. These experiments were performed to showcase:
  1. (a)
    Comparing our methodology to out-of-the-box solutions for article embeddings.
  2. (b)
    Evaluating if our semi-supervised approach yields an agnostic improvement to the generated articles embedding, relative to the given basic article embedding representation.
To perform this comparison, we also included in Table 3 the results of the basic spaCy model and the semi-supervised infused spaCy model. Additionally, to determine the significant improvement of our approach, we included Figure7, which evaluated the performance of each model embedding for all the experiments of common event attribution.
(a) Precision-Recall AUC Comparison
(b) ROC AUC Comparison
@@ 5.4 Statistical significance evaluation
We perform statistical analysis to verify the existence of significant differences among the evaluated methods. As suggested by [72], we use the Friedman test [68] on the results of the aforementioned approaches to identify which obtained better results regarding each designed task and metric across all datasets.
The tests were used with a significance level of 5%. Specifically, the rank results established by the statistical method provide the criterion for evaluating the approaches, considering that the higher the value of the calculated rank, the better the result will be in this approach. In our case, the null hypothesis was rejected in each context, as can be seen in Table 5. The rejection of the null hypothesis of the Friedman test indicates the existence of statistical differences between the methods but does not define which of them are statistically different. For that, the Nemenyi post-hoc test [69] was applied to make multiple comparisons using the values of the percentages of means for each approach evaluated in all the executed datasets, as can be seen in Table 6. The results of the statistical analysis are presented in Appendix A.4.
Therefore, in all cases, there were significant differences between the baseline SIF approach and the Concatenated approach. By the Precision-Recall metric, there were also significant differences between our proposed SiameseNet and the Concatenated approach. Regarding the spaCy model, there were significant differences between the spaCy model performance compared to all methods that relied on our suggested time-based entity embedding or semi-supervised event approach.
@@ 6 Discussion
Upon analyzing the results presented in the previous section, we can conclude the following: First, the utilization of entities, themes, and events to construct news embeddings has shown promising results, as reflected in Tables 1 and 3. Even though the task of shared event detection between pairs of articles in a given timeframe is challenging, as discussed in the Introduction section, all proposed methods achieved extremely high ROC AUC scores, all above 0.95, while also providing impressive results concerning the more challenging metric of PR AUC.
@@
Second, utilizing the semi-supervised approach, which considers cross-document latent information given the tagged events, contributes to our method performance. From Figure 6, it is evident that the semi-supervised approach achieves significant improvement compared to the SIF baseline by the ROC AUC metric on all tasks. Additionally, from analyzing Figure 8, we see that even though there is a high correlation between the methods’ performance, a variance in performance exists. Namely, we found some cases where the baseline performed better than the semi-supervised approach. We attribute this behavior to two reasons: (a) we found a moderate correlation of 0.369 for each dataset size and the gap in performance between the SIF and semi-supervised approach; more data led to a bigger gap in favor of the Siamese network, (b) a lack of adaptation of the Siamese network to the given dataset or task led to variances in its results.
@@
Third, our method, which generates news article embeddings, produced embeddings for articles and entities on a large scale of news data without GPU. Using the resources described in Appendix A, we ran every step of the pipeline and every experiment on all the generated datasets in parallel. For each entity and theme we monitored, over 320,000 overall (see Table 2), we generated topical embeddings for every period they appeared in from 07/2015 to 01/2021. We have considered all articles from dozens of monitored media sources and their affiliates, as presented in Figure 4, resulting in a dataset of over 850,000 articles. This enables researchers with fewer resources to investigate large-scale online news using this method.
@@
Fourth, we constructed topical monthly entity embeddings for over 320,000 distinct entities and themes. The pooling mechanism that controlled the training set of the GloVe algorithm served as a critical point in this research. This mechanism not only allowed us to control the resource utilization of our suggested method but also fine-tuned the representation of each entity to a recent time frame, such as its article’s publication date. In Table 4, we further explain and demonstrate real-world examples of the importance of topical embeddings. We plan to share these topical entity embeddings with the research community. These topical embeddings could serve other studies examining how relationships between entities change over time.
Fifth, as presented in the previous section, the experimental results demonstrate the superiority of our proposed method compared to the baseline method SIF and the biggest CPU spaCy method. This can be observed in Table 3. We detected only three instances out of 66 datasets and two tasks in which the SIF approach surpassed our technique. The spaCy model and the Semi-Supervised infused spaCy embeddings fell significantly short compared to all other approaches in all of the designed experiments and measured metrics. Additionally, Figure 7 shows that our Semi-Supervised approach significantly improves the spaCy embeddings for news articles for the task of common event attribution.
@@
Sixth, we observed that in the cases in which better performance is yielded by the Siamese network or SIF Methods compared to the Concatenated approach, there was a high gap between the SIF approach and the Siamese approach, leading to the failure of the concatenated approach to successfully account for additional information by utilizing both approaches and resulting in lesser performance than taking the best approach given the dataset.
@@
Lastly, the method presented in this research has some limitations worth noting: First, it relies on the historical existence of event tagging for news articles, which might not always be available or reliable enough. Second, as suggested by the results, our methodology is relatively static, and the pipeline should preferably be manually tuned to provide adaptive embeddings to the required task; these concerns both in the semi-supervised training process and in the feature selection phase.
@@ 7 Conclusions and Future Work
This research has presented a novel approach for generating event-based embeddings for news articles, addressing the limitations of existing methods that rely solely on the textual content of documents, disregard runtime and memory costs, ignore the contextual time of article publication date, or do not rely on cross-document information. We do so by leveraging the historical context of entities, themes, and their associations with specific events. Our proposed method aimed to capture the latent relationships between these elements, ultimately enhancing the representation of news articles in a vector space.
We demonstrated our method’s ability to produce article embeddings that consider the article’s published time and represent the latent event information that appears in them, all while using relatively low resources on large-scale datasets. We also showed the method’s ability to surpass and improve existing solutions.
This research offers multiple directions for possible future work: The first direction is investigating the interpretability and explainability of the generated embeddings, which could be a valuable research direction. Understanding the factors that contribute to the embeddings’ representations and their relationships with events and entities could lead to more transparent and trustworthy models and facilitate the incorporation of human expertise and domain knowledge into the embedding generation process.
This also paves the way for analyzing the relationships between certain entities and tense or biased media coverage. In the future, we may apply a holistic approach to event clustering based on the methods proposed in this work to facilitate possible research in media slant or political news.
@@ A.3 Embedding Generation Notes
Entity Entity Neighbor GloVe GloVe+Pool Mechanism
@@
Fifth RS Alexander_Litvinenko